OpenStudy (henrietepurina):

Algebra 1 question Medal + Fan

OpenStudy (henrietepurina):

Using the equation below as a model, fill in numbers in the place of a and b to create a rational equation that has an extraneous solution. x plus a over ax = b over x Part 1. Show all work to solve for x in the equation and check the solution. Part 2. Explain how to identify the extraneous solution and what it means.

OpenStudy (henrietepurina):

Can you help @surjithayer ?

OpenStudy (anonymous):

is this your statement? \[x+\frac{ a }{ ax }=\frac{ b }{ x }\]

OpenStudy (henrietepurina):

not quite... ill post a photo

OpenStudy (henrietepurina):

If I choose a=1 and b=2, would it still have a extraneous solution?

jimthompson5910 (jim_thompson5910):

Are you able to solve for x without having to replace 'a' and 'b' with numbers?

OpenStudy (henrietepurina):

Im not supposed to, but I guess I probably could if needed...

jimthompson5910 (jim_thompson5910):

That's what I recommend. Tell me what you get (you should get 2 solutions).

OpenStudy (henrietepurina):

hmm.... I got b=((a+x)/(b)) and a=((−x)/(−b+1))

jimthompson5910 (jim_thompson5910):

you need to solve for x

OpenStudy (henrietepurina):

ok then, ax+x^2=abx and x=ab-a

jimthompson5910 (jim_thompson5910):

I was thinking more along like this... (x+a)/(ax) = b/x x(x+a) = ax*b ... cross multiply x(x+a) - ax*b = 0 x(x+a - ab) = 0 ... factor x = 0 or x+a - ab = 0 x = 0 or x+a = ab x = 0 or x = ab-a

jimthompson5910 (jim_thompson5910):

Notice how no matter what 'a' and 'b' are, the possible solution x = 0 pops up. However, we must check every possible solution back into the original equation. (x+a)/(ax) = b/x (0+a)/(a*0) = b/0 ... and we run into a division by zero error here so x = 0 is an extraneous solution

OpenStudy (anonymous):

\[x^2+ax=abx\] \[x^2+ax-abx=0\] \[x^2-(ab-a)x=0\] \[x[x-a(b-1)]=0\] either x=0 but \[x \neq 0\] therefore x is an extraneous root,a & b can have any value. or x=a(b-1)

jimthompson5910 (jim_thompson5910):

x = ab-a is the only possible solution left, but it cannot be zero or else you'll run into the same issues above x = ab-a x = a(b-1) if x = 0, then a(b-1) = 0 which leads to a = 0 or b-1 = 0 --> b = 1 So if a = 0 or b = 1 (or both), then x = ab-a will be zero. That means we will have another extraneous solution. If 'a' is not zero, and b is not 1, then you'll have no division by zero issues to worry about.

OpenStudy (henrietepurina):

Ok.... data overload :0.... can you do a short recap of what you just said?

jimthompson5910 (jim_thompson5910):

I solved for x in general (in terms of 'a' and 'b'). When I did that I got these 2 possible solutions x = 0 or x = ab - a The first solution, x = 0, doesn't work in the original equation. So it's extraneous

OpenStudy (anonymous):

here for x to be extraneous root \[a \neq0,b \neq 1\]

jimthompson5910 (jim_thompson5910):

the only other solution x = ab-a can be any number but zero to make sure ab-a is nonzero, we must make sure that 'a' is not zero and b is not 1

OpenStudy (henrietepurina):

thank you...

OpenStudy (henrietepurina):

Are we done with part A?

jimthompson5910 (jim_thompson5910):

well I haven't done part A or part B. I just solved it in general to show you that no matter what numbers you pick for 'a' and 'b', you'll always get at least 1 extraneous solution

jimthompson5910 (jim_thompson5910):

you said "If I choose a=1 and b=2, would it still have a extraneous solution?" the answer is "yes, because x = 0 is always one of the possible solutions, so x = 0 is extraneous" and you would show why x = 0 is extraneous

OpenStudy (henrietepurina):

Alright 'o, thanks for clarifying that :D

jimthompson5910 (jim_thompson5910):

with a = 1 and b = 2, this means x = ab-a x = 1*2-1 ... plug in a = 1 and b = 2 x = 2-1 x = 1 So if a = 1 and b = 2, then x = 1 is the solution (turns out the only solution)

jimthompson5910 (jim_thompson5910):

or you could pick a = 1 and b = 2, plug them into the original equation, then solve for x

OpenStudy (henrietepurina):

Thanks a ton, now were off to B.... right?

jimthompson5910 (jim_thompson5910):

yeah you just need to explain how to ID the extraneous solutions

OpenStudy (henrietepurina):

ID... like identify?

jimthompson5910 (jim_thompson5910):

yeah, sry I was being lazy with typing lol

OpenStudy (henrietepurina):

no no thats ok.... would you just simply say: Once you solve the equation you're given, take the solutions and enter them into the original equation. If the answer comes out to be equal to what its supposed to be equal to, then the solution is valid. If the answer is not equal to what it was supposed to be equal to, then the solution is extraneous.

jimthompson5910 (jim_thompson5910):

yes, also tack on "if plugging a possible solution back into the original problem causes a division by zero error, then that possible solution isn't really a true solution at all. It's extraneous" or something like that It turns out that it all has to do with domain. If you get a possible solution that is not in the domain, then it's not a true solution. This solution is extraneous.

OpenStudy (henrietepurina):

OK... did that :) Time for part 2

jimthompson5910 (jim_thompson5910):

I thought that was part 2?

jimthompson5910 (jim_thompson5910):

part 2 is " Explain how to identify the extraneous solution and what it means."

OpenStudy (henrietepurina):

OOOPS OMG I MESSED UP THANKS A TONGAZILION :D

OpenStudy (henrietepurina):

Can you help me on another question?

jimthompson5910 (jim_thompson5910):

sure, what is it?

OpenStudy (henrietepurina):

Create a quadratic polynomial function f(x) and a linear binomial in the form (x − a). Part 1. Show all work using long division to divide your polynomial by the binomial. Part 2. Show all work to evaluate f(a) using the function you created. Part 3. Use complete sentences to explain how the remainder theorem is used to determine whether your linear binomial is a factor of your polynomial function

jimthompson5910 (jim_thompson5910):

alright, go ahead and post what you have so far, I'll brb in a few min

jimthompson5910 (jim_thompson5910):

ok back